# Conceptualization, Methodology, Software, Validation, Formal analysis, Data Curation, Investigation, Writing – original draft, Writing - Review & Editing, Visualization

### v1

@@ 4.1 Pre-training
where $sim(\mathbf{E}_{m},\mathbf{E}_{mt})=\mathbf{E}_{m}\cdot\mathbf{E}_{mt}/(\|\mathbf{E}_{m}\|\|\mathbf{E}_{mt}\|)$ is the cosine similarity and $\tau$ is a temperature parameter which can be seen as weight for the most similar negative pair.
@@
For molecular substructures, existing work demonstrates that the properties of drug molecules partially depend on specific substructures, and the treatment of a patient’s disease also relies on the functions of certain substructures. Our goal is to reconstruct the 3D structures of substructures to make them closer to the actual molecular structures. We first use the BRICS [33] method to decompose drug molecules into substructures, and then complete the 3D information of the substructures using the API provided by RDKit. We decompose all drug molecules to obtain a substructure set $\mathcal{S}$, and finally obtain the embedding matrices of the two different modalities of the structures $\mathbf{E}_{s},\mathbf{E}_{st}\in\mathbb{R}^{|\mathcal{S}|\times h}$, where each row represents the embedding of a specific substructure.
4.2 Molecular Multi-Distillation
@@
We encode the current diagnoses, surgeries, and medications as the health environment of the patient at the current time sequence. We first extract the sets of procedures, diagnoses, and medications at each time point from the visit $v(t-1)$: $d(t-1)=\{d_{i},d_{j},\ldots\}$, $p(t-1)=\{p_{i},p_{j},\ldots\}$, and $m(t-1)=\{m_{i},m_{j},\ldots\}$, all of which are multi-hot vectors mentioned in our problem definition. We define three learnable embedding matrices, $\mathbf{E}_{d}$, $\mathbf{E}_{p}$, and $\mathbf{E}_{me}$, corresponding to diagnoses, surgeries, and medications, where dim is the embedding size. We generate $\mathbf{e}_{d_{i}}$, $\mathbf{e}_{p_{i}}$, and $\mathbf{e}_{m_{i}}$ by mapping diagnoses $d_{i}$, surgeries $p_{i}$, and medications $m_{i}$ to the embedding space, and use them as entity-level representations shared globally for downstream tasks. We obtain the set representations $e_{d}(t)$, $e_{p}(t)$, and $e_{m}(t)$ by mapping all diagnoses, surgeries, and medications through their corresponding embedding matrices.
$e_{d}(t-1)=d(t-1)E_{d},$ (6)
$e_{p}(t-1)=p(t-1)E_{p},$ (7)
$e_{m}(t-1)=m(t-1)E_{m}.$ (8)
Finally, we concatenate the three vectors of equal length, resulting in a single vector with three times the original length, forming the visit-level representation at time $t-1$, $e_{v}^{\prime}(t)\in R^{1\times 3dim}$.
$e_{v}^{\prime}(t-1)=\mathrm{CONCAT}(e_{d}(t-1),e_{p}(t-1),e_{m}(t-1)).$ (9)
Then, a multi-layer perceptron(MLP) converts the dimension of the vector $e_{v}^{\prime}(t-1)$ to $|\mathcal{S}|$, the total number of molecular substructures, and a sigmoid function determines the degree to which each substructure is relevant at the current time.
$e_{s}^{\prime}=\mathrm{MLP}(e_{v}^{\prime}(t-1))$ (10)
@@
We use the molecular structure encoder trained during the pretraining phase to encode molecular structures and their substructures, obtaining $\mathbf{E}^{\prime}_{s}$ and $\mathbf{E}^{\prime}_{m}$. For substructures, since the function of molecules is realized through a set of substructures, we need to model the interaction information of substructures. Therefore, we design a substructures self-attention(SSA) module to achieve this goal. We use the embedding matrix $\mathbf{E}_{s}$ of all substructures as input to establish an attention mechanism between each substructure, and finally output an embedding matrix of equal size $\mathbf{E}^{*}_{s}\in\mathbb{R}^{|\mathcal{S}|\times h}$. The self-attention layer is defined as follows:
$\mathrm{H=LayerNorm(E_{s}+\mathcal{Attn}_{self}(E_{s})),}$ (11)
$\mathrm{E_{s}^{*}=SSA(E_{s})=LayerNorm(H+MLP(H)),}$ (12)
$\mathcal{Attn}_{self}$ denotes the self-attention mechanism defined by the following formula:
$\mathcal{Attn}_{self}(\mathrm{X})=\mathrm{Softmax}\left(\frac{\mathcal{Q}_{x}\mathcal{K}_{x}^{\top}}{\sqrt{d_{k}}}\right)\mathcal{V}_{x},$ (13)
where $d_{k}$ is the dimension of the key embeddings, and $\mathbf{Q}_{x}$, $\mathbf{K}_{x}$, and $\mathbf{V}_{x}$ are the query, key, and value matrices of the input $\mathbf{X}$.
We perform attention aggregation on the overall molecular representation $\mathbf{E}^{\prime}_{m}$ and the substructure interaction representation $\mathbf{E}^{*}_{s}$, using a mask matrix to rescale the attention matrix when calculating the attention coefficients.
$\mathcal{Attn}_{mol-sub}(\mathrm{X,Y})=\mathrm{Softmax}\left(\frac{\mathcal{Q}_{x}\mathcal{K}_{y}^{\top}}{\sqrt{d_{m}in}}\right),$ (14)
$\mathcal{Attn}_{mol-sub}(\mathrm{X,Y})=\mathrm{mask}(\mathcal{Attn}_{mol-sub},-(1\ll 32))$ (15)
@@
$\forall s\in S_{k},\tilde{\mathbf{y}}_{s}^{(i)}=\mathbf{e}_{s}^{\prime(i)}\mathbf{y}_{s},$ (16)
where $\mathbf{e}_{s}^{\prime(i)}$ is the element of $\mathbf{e}_{s}^{\prime}$ corresponding to the substructure $s$. Finally, we obtain the perception representation of the substructure at the current time sequence through the attention matrix.
$\ddot{\mathbf{y}}_{k}^{(i)}=\sum_{s\in S_{k}}\mathcal{Attn}_{mol-sub}\tilde{\mathbf{y}}_{s}^{(i)}.$ (17)
@@
$e_{d}(t)=d(t)E_{d},e_{p}(t)=p(t)E_{p},$ (18)
$e_{v}^{\prime}(t)=\mathrm{CONCAT}(e_{d}(t),e_{p}(t)),$ (19)
$\mathbf{W}_{m}=\sigma(\mathrm{MLP}(e_{v}^{\prime}(t))).$ (20)
@@
4.3 Fusion of Drug Evolution Trajectories
The purpose of this stage is to integrate the drug representations obtained for the patient at each time point. This representation results from the interaction of deep and shallow knowledge and subsequent bias correction, ultimately yielding the recommendation probabilities for all drugs.
By embedding the multiple drug representations generated in the previous stage $\{\hat{\mathbf{m}}_{v_{1}},\hat{\mathbf{m}}_{v_{2}},\ldots,\hat{\mathbf{m}}_{v_{t-1}}\}$ and inputting them into a gated recurrent unit $\text{GRU}(\cdot)\in\mathbb{R}^{|\mathcal{M}|}$ to model the sequential relationships, we then process the outputs through a multilayer perceptron
@@ 4.4 Model Training
Medication recommendation is essentially a multi-label binary classification task. Therefore, we use the binary cross-entropy loss function $L_{bce}$ and the multi-label margin loss function $L_{multi}$. Additionally, due to the specific nature of drug recommendation, we use the drug DDI loss function $L_{ddi}$ [33], which calculates the probability of DDI occurrence by identifying drug pairs with potential DDI risk in the drug combination. The definitions of the three loss functions are as follows:
@@
Although DDI indicates the potential risk of drug combinations, considering medical practice, certain adverse reactions are acceptable. Pursuing the lowest DDI rate without considering efficacy does not lead to the best drug combination. Therefore, we must set a threshold for DDI; if this threshold is exceeded, the weight of DDI increases. When using the composite loss function, we adopt the same research methodology as before [33] to balance the accuracy and safety of the drug combinations.
@@ 5.1.3 Sampling Approach
Due to the limited availability of publicly accessible EHR data, we adopt bootstrapping sampling in this phase, following the approach recommended in [32]. This technique is particularly effective in scenarios where sample sizes are small, as detailed in [19] and [8].
@@ 5.3 Evaluation Metrics
We delve into the performance evaluation of our method using four principal metrics: Jaccard, DDI rate, F1, and PRAUC. Here’s an in-depth explanation of each metric and its application in our study.
@@
where $\hat{m}_{i}$ represents the multi-hot vector of the predicted outcome, $m_{i}$ represents the real label, $Jaccard(t)$ represents the evaluation result at visit $t$, and $N_{h}$ represents the total number of visits for patient $h$.
@@
where $N_{h}$ denotes the total number of visits for patient $h$, $m(t)$ and $\hat{m}(t)$ denote the real and predicted multi-labels at visit $t$, $m_{j}(t)$ denotes the $j^{th}$ entry of $m(t)$, $a^{ddi}$ is the prior DDI relation matrix, and $1$ is an indicator function that returns 1 when $a^{ddi}=1$, otherwise 0.
@@
Avg. # of Drugs (Average number of drugs) measures the average number of drugs in each recommendation result. A higher value indicates that each recommendation contains more drugs, which may increase the complexity of clinical treatment and the risk of adverse reactions. Conversely, a lower value suggests that the drug combinations may be safer and help minimize unnecessary drug use. It is important to note that this metric is for reference only and the size of the combination should not be used as a strict evaluation standard.
@@
where $N_{h}$ represents the total number of visits for patient $h$ and $|\hat{M}(i)|)$ denotes the number of predicted medications in visit $i$ of patient $h$.
@@ 5.5 Performance Comparison
Our model re-calibrates the representation of molecules and their substructures by introducing the 3D modality of molecules, increasing the dimensionality of molecular representation, addressing real structural confusion issues, and designing a substructure distillation mechanism at the access level. This adjusts the interaction mechanism between molecular knowledge and patients’ physical conditions, enhancing model rationality and accuracy while only slightly increasing the DDI rate.
@@ 6.2 Efficiency Analysis
In the subsequent substructure interaction distillation stage, we use three embedding layers, a GRU layer, a molecular structure and its substructure graph neural network along with the corresponding attention mechanism, and multiple fully connected layers. The time complexity of the embedding layers is $O(3\cdot n\cdot d)$, where $n$ is the length of the input sequence and $d$ is the size of the embedding dimension. The time complexity of the GRU layer is $O(t\cdot h^{2})$, where $t$ is the sequence length and $h$ is the size of the hidden layer. The time complexity of the GNN layer is $O(2\cdot(N+E)^{4})$, where $N$ is the number of nodes and $E$ is the number of edges. The time complexity of the attention mechanism is $O(2\cdot L^{2}\cdot D)$, where $L$ is the length of the input sequence and $D$ is the corresponding dimension. The time complexity of multiple fully connected layers is denoted as $O(\Sigma(n_{i}\cdot n_{o}))$, where $n_{i}$ is the number of input nodes and $n_{o}$ is the number of output nodes. Finally, the overall time complexity is:
@@ 6.3 Ablation Study
BiMoRec $w/o$ A 0.5356 0.0697 0.6887 0.7823 20.4086 0.4874 0.0673 0.6382 0.7318 14.8917
BiMoRec $w/o$ B 0.5373 0.0727 0.6906 0.7787 22.0278 0.4836 0.0649 0.6346 0.7244 14.8244
BiMoRec $w/o$ A+B 0.5357 0.0672 0.6888 0.7815 22.1939 0.4798 0.0622 0.6307 0.7209 14.7726
BiMoRec 0.5526 0.0684 0.7033 0.7955 22.4693 0.4912 0.0635 0.6419 0.7324 14.8019
@@
BiMoRec $w/o$ B: We eliminate the Substructure-Patient Interaction Distillation module in this variant, removing the interaction between substructures and a single patient visit. Instead, we rely solely on the patient’s latest physical condition to interact with the single visit representation, and finally obtain the weights of the medication substructures.
BiMoRec $w/o$ A+B: We simultaneously remove both the pre-training module and the Substructure-Patient Interaction Distillation module. This variant neither integrates the dual modalities of molecular representation nor interacts with the patient’s single representation, relying entirely on the relationships between medications, diseases, and surgeries in patient visits to make recommendations.

### v2

@@ 4.1 Pre-training
where $sim(\mathbf{E}_{m},\mathbf{E}_{n})=\mathbf{E}_{m}\cdot\mathbf{E}_{n}/(\|\mathbf{E}_{m}\|\|\mathbf{E}_{n}\|)$ is the cosine similarity and $\tau$ is a temperature parameter which can be seen as weight for the most similar negative pair.
@@
For molecular substructures, existing work demonstrates that the properties of drug molecules partially depend on specific substructures, and the treatment of a patient’s disease also relies on the functions of certain substructures. Our goal is to reconstruct the 3D structures of substructures to make them closer to the actual molecular structures. We first use the BRICS [43] method to decompose drug molecules into substructures, and then complete the 3D information of the substructures using the API provided by RDKit [18]. We decompose all drug molecules to obtain a substructure set $\mathcal{S}$, and finally obtain the embedding matrices of the two different modalities of the substructures $\mathbf{E}_{s},\mathbf{E}_{st}\in\mathbb{R}^{|\mathcal{S}|\times h}$, where each row represents the embedding of a specific substructure.
We apply this loss function to molecules and their substructures, inputting $\mathbf{E}_{m}$ and $\mathbf{E}_{mt}$ into the above formula to obtain $\mathcal{L}_{mol}$, and inputting $\mathbf{E}_{s}$ and $\mathbf{E}_{st}$ into the above formula to obtain $\mathcal{L}_{submol}$. Finally, we sum them to obtain the total loss function.
$\mathcal{L}_{total}=\mathcal{L}(\mathbf{E}_{m},\mathbf{E}_{mt})+\mathcal{L}(\mathbf{E}_{s},\mathbf{E}_{st})$ (10)
4.2 Molecular Multi-Enhancement
@@
We encode the current diagnoses, surgeries, and medications as the health environment of the patient at the current time sequence. We first extract the sets of procedures, diagnoses, and medications at each time point from the visit $v(t-1)$: $d(t-1)=\{d_{i},d_{j},\ldots\}$, $p(t-1)=\{p_{i},p_{j},\ldots\}$, and $m(t-1)=\{m_{i},m_{j},\ldots\}$, all of which are multi-hot vectors mentioned in our problem definition. We define three learnable embedding matrices, $\mathbf{E}_{d}$, $\mathbf{E}_{p}$, and $\mathbf{E}_{m}$, corresponding to diagnoses, surgeries, and medications. We generate $\mathbf{e}_{d_{i}}$, $\mathbf{e}_{p_{i}}$, and $\mathbf{e}_{m_{i}}$ by mapping diagnoses $d_{i}$, surgeries $p_{i}$, and medications $m_{i}$ to the embedding space, and use them as entity-level representations shared globally for downstream tasks. We obtain the set representations $\mathbf{e}_{d}(t)$, $\mathbf{e}_{p}(t)$, and $\mathbf{e}_{m}(t)$ by mapping all diagnoses, surgeries, and medications through their corresponding embedding matrices.
$\mathbf{e}_{d}(t-1)=d(t-1)\mathbf{E}_{d},$ (11)
$\mathbf{e}_{p}(t-1)=p(t-1)\mathbf{E}_{p},$ (12)
$\mathbf{e}_{m}(t-1)=m(t-1)\mathbf{E}_{m}.$ (13)
Finally, we concatenate the three vectors of equal length, resulting in a single vector with three times the original length, forming the visit-level representation at time $t-1$, $\mathbf{e}_{v}^{\prime}(t)\in R^{1\times 3h}$, where $h$ is the embedding size.
$\mathbf{e}_{v}^{\prime}(t-1)=\mathrm{CONCAT}(\mathbf{e}_{d}(t-1),\mathbf{e}_{p}(t-1),\mathbf{e}_{m}(t-1)).$ (14)
Then, a multi-layer perceptron(MLP) converts the dimension of the vector $\mathbf{e}_{v}^{\prime}(t-1)$ to $|\mathcal{S}|$, the total number of molecular substructures, and a sigmoid function determines the degree to which each substructure is relevant at the current time.
$\mathbf{e}_{s}^{\prime}=\mathrm{MLP}(\mathbf{e}_{v}^{\prime}(t-1))$ (15)
@@
We use the molecular structure encoder trained during the pretraining phase to encode molecular structures and their substructures, obtaining $\mathbf{E}^{\prime}_{s}$ and $\mathbf{E}^{\prime}_{m}$. For substructures, since the function of molecules is realized through a set of substructures, we need to model the interaction information of substructures. Therefore, we design a substructures self-attention(SSA) module [21] to achieve this goal. We use the embedding matrix $\mathbf{E}^{\prime}_{s}$ of all substructures as input to establish an attention mechanism between each substructure, and finally output an embedding matrix of equal size $\mathbf{E}^{*}_{s}\in\mathbb{R}^{|\mathcal{S}|\times h}$. The self-attention layer is defined as follows:
$\mathrm{H=LayerNorm(\mathbf{E}^{\prime}_{s}+\mathbf{A}_{self}(\mathbf{E}^{\prime}_{s})),}$ (16)
$\mathrm{\mathbf{E}^{*}_{s}=SSA(\mathbf{E}^{\prime}_{s})=LayerNorm(H+MLP(H)),}$ (17)
$\mathcal{Attn}_{self}$ denotes the self-attention mechanism [34] defined by the following formula:
$\mathbf{A}_{self}(\mathrm{X})=\mathrm{Softmax}\left(\frac{\mathbf{Q}\mathbf{K}^{\top}}{\sqrt{d}}\right)\mathbf{V},$ (18)
where $d$ is the dimension of the key embeddings, and $\mathbf{Q}$, $\mathbf{K}$, and $\mathbf{V}$ are the query, key, and value matrices of the input $\mathbf{X}$.
For molecular structures, we need to select from the embedding table $\mathbf{E}^{\prime}_{m}$ based on the patient’s visit data and perform molecular enhancement. We input $\mathbf{E}^{\prime}_{m}$ into the global selection enhancement module (GSE) to obtain the enhanced result $\mathbf{E}^{*}_{m}$:
$\mathbf{E}^{*}_{m}=\text{GSE}(\text{SSA}(\mathbf{E}^{\prime}_{m}[i]\mid i\in m(t-1)))$ (19)
We perform attention aggregation on the overall molecular representation $\mathbf{E}^{*}_{m}$ and the substructure interaction representation $\mathbf{E}^{*}_{s}$, using a mask matrix to rescale the attention matrix when calculating the attention coefficients.
$\mathbf{A}_{mol-sub}(\mathrm{X,Y})=\mathrm{Softmax}\left(\frac{\mathbf{Q}_{x}\mathbf{K}_{y}^{\top}}{\sqrt{d_{min}}}\right),$ (20)
$\mathbf{A}_{mol-sub}(\mathrm{X,Y})=\mathrm{mask}(\mathbf{A}_{mol-sub},-(1\ll 32))$ (21)
@@
$\forall s\in S,\tilde{\mathbf{y}}_{s}^{(i)}=\mathbf{e}_{s}^{\prime(i)}\mathbf{y}_{s},$ (22)
where $\mathbf{e}_{s}^{\prime(i)}$ is the element of $\mathbf{e}_{s}^{\prime}$ corresponding to the substructure $s$. We obtain the perception representation of the substructure at the current time sequence through the attention matrix.
$\mathbf{y}_{m}^{(i)}=\sum_{s\in S}\mathbf{A}_{mol-sub}\tilde{\mathbf{y}}_{s}^{(i)}.$ (23)
@@
$\mathbf{e}_{d}(t)=d(t)\mathbf{E}_{d},\mathbf{e}_{p}(t)=p(t)\mathbf{E}_{p},$ (24)
$\mathbf{e}_{v}^{\prime}(t)=\mathrm{CONCAT}(\mathbf{e}_{d}(t),\mathbf{e}_{p}(t)),$ (25)
$\mathbf{W}_{m}=\sigma(\mathrm{MLP}(\mathbf{e}_{v}^{\prime}(t))).$ (26)
@@
4.3 Fusion of Medication Evolution Trajectories
The purpose of this stage is to integrate the medication representations obtained for the patient at each time point. This representation results from multiple dynamic enhancements of substructures, ultimately yielding the recommendation probabilities for all medications.
By embedding the multiple medication representations generated in the previous stage $\{\hat{\mathbf{m}}_{v_{1}},\hat{\mathbf{m}}_{v_{2}},\ldots,\hat{\mathbf{m}}_{v_{t-1}}\}$ and inputting them into a gated recurrent unit $\text{GRU}(\cdot)\in\mathbb{R}^{|\mathcal{M}|}$ to model the sequential relationships, $\{v_{1},v_{2},\ldots,v_{t-1}\}$ are the first $t-1$ patient visits, we then process the outputs through a multilayer perceptron
@@ 4.4 Model Training
Medication recommendation is essentially a multi-label binary classification task. Therefore, we use the binary cross-entropy loss function $L_{bce}$ and the multi-label margin loss function $L_{multi}$. Additionally, due to the specific nature of medication recommendation, we use the medication DDI loss function $L_{ddi}$ [43], which calculates the probability of DDI occurrence by identifying medication pairs with potential DDI risk in the medication combination. The definitions of the three loss functions are as follows:
@@
Although DDI indicates the potential risk of drug combinations, considering medical practice, certain adverse reactions are acceptable. Pursuing the lowest DDI rate without considering efficacy does not lead to the best drug combination. Therefore, we must set a threshold for DDI; if this threshold is exceeded, the weight of DDI increases. When using the composite loss function, we adopt the same research methodology as before [43, 9] to balance the accuracy and safety of the medication combinations.
@@ 5.1.3 Sampling Approach
Given the limited availability of publicly accessible EHR data, we employ bootstrapping sampling during this phase, following the methodology suggested in [42]. As discussed in [28] and [8], this technique proves particularly effective in scenarios with small sample sizes.
@@ 5.2 Datasets
We obtain the molecular SDF files from the PubChem database [17], and the 3D data for substructures is completed using functions from the RDKit library.
@@ 5.3 Evaluation Metrics
We use four primary metrics, Jaccard, DDI rate, F1, and PRAUC, to evaluate the performance of our method. The specific application of each metric in our study is explained below.
@@
where $\hat{m}_{i}(t)$ represents the multi-hot vector of the predicted outcome, $m_{i}$ represents the real label, $Jaccard(t)$ represents the evaluation result at visit $t$, and $N_{h}$ represents the total number of visits for patient $h$.
@@
where $N_{h}$ denotes the total number of visits for patient $h$, $m_{j}(t)$ denotes the $j^{th}$ entry of $m(t)$, $m(t)$ and $\hat{m}(t)$ denote the real and predicted multi-labels at visit $t$, $a^{ddi}$ is the prior DDI relation matrix, and $1$ is an indicator function that returns 1 when $a^{ddi}=1$, otherwise 0.
@@
Avg. # of Drugs (Average number of drugs) measures the mean number of medications in each recommendation result. A higher value indicates that each recommendation includes more medications, potentially increasing the complexity of clinical treatment and the risk of adverse reactions. Conversely, a lower value suggests that the medication combinations may be safer and help minimize unnecessary medication use. It is important to note that this metric serves as a reference only, and the size of the combination should not be used as a strict evaluation standard.
@@
where $N_{h}$ represents the total number of visits for patient $h$ and $|\hat{M}(i)|$ denotes the number of predicted medications in visit $i$ of patient $h$.
@@ 5.5 Performance Comparison
Our model introduces the 3D modality of molecules, recalibrating the representations of molecules and their substructures, thus increasing the dimensionality of molecular representations and addressing the existing issue of structural confusion. We design a visit-level substructure enhancement mechanism, adjusting the interaction mechanism between molecular knowledge and the patient’s health condition, enhancing the model’s rationality and accuracy. This approach only slightly increases the DDI rate, ensuring a balance between accuracy and safety.
@@ 6.2 Efficiency Analysis
In the subsequent substructure multi-enhancement stage, we use three embedding layers, a GRU layer, a molecular structure and its substructure graph neural network along with the corresponding attention mechanism, and multiple fully connected layers. The time complexity of the embedding layers is $O(3\cdot n\cdot d)$, where $n$ is the length of the input sequence and $d$ is the size of the embedding dimension. The time complexity of the GRU layer is $O(t\cdot h^{2})$, where $t$ is the sequence length and $h$ is the size of the hidden layer. The time complexity of the GNN layer is $O(2\cdot(N+E)^{4})$, where $N$ is the number of nodes and $E$ is the number of edges. The time complexity of the attention mechanism is $O(2\cdot L^{2}\cdot D)$, where $L$ is the length of the input sequence and $D$ is the corresponding dimension. The time complexity of multiple fully connected layers is denoted as $O(\Sigma(n_{i}\cdot n_{o}))$, where $n_{i}$ is the number of input nodes and $n_{o}$ is the number of output nodes. Finally, the overall time complexity is:
@@ 6.3 Ablation Study
BiMoRec $w/o$ A 0.5356 0.0697 0.6887 0.7823 20.4086 0.4832 0.0704 0.6348 0.7268 13.5401
BiMoRec $w/o$ B 0.5347 0.0735 0.6882 0.7752 21.8691 0.4800 0.0698 0.6320 0.7231 15.5599
BiMoRec $w/o$ A+B 0.5335 0.0747 0.6873 0.7782 22.7640 0.4778 0.0696 0.6301 0.7212 15.9020
BiMoRec 0.5422 0.0707 0.6944 0.7858 20.3989 0.4855 0.0696 0.6369 0.7315 13.7313
@@
BiMoRec $w/o$ B: We eliminate the molecular multi-enhancement module in this variant, removing the interaction between substructures and a single patient visit. Instead, we rely solely on the patient’s latest physical condition to interact with the single visit representation, and finally obtain the weights of the medication substructures.
BiMoRec $w/o$ A+B: We simultaneously remove both the pre-training module and the molecular multi-enhancement module. This variant neither integrates the dual modalities of molecular representation nor interacts with the patient’s single representation, relying entirely on the relationships between medications, diseases, and surgeries in patient visits to make recommendations.